# Can Large Language Models Replace Data Scientists in Clinical Research?

### v1

@@ Abstract
Data science plays a critical role in clinical research, but it requires professionals with expertise in coding and medical data analysis. Large language models (LLMs) have shown great potential in supporting medical tasks and performing well in general coding tests. However, these tests do not assess LLMs’ ability to handle data science tasks in medicine, nor do they explore their practical utility in clinical research. To address this, we developed a dataset consisting of 293 real-world data science coding tasks, based on 39 published clinical studies, covering 128 tasks in Python and 165 tasks in R. This dataset simulates realistic clinical research scenarios using patient data. Our findings reveal that cutting-edge LLMs struggle to generate perfect solutions, frequently failing to follow input instructions, understand target data, and adhere to standard analysis practices. Consequently, LLMs are not yet ready to fully automate data science tasks. We benchmarked advanced adaptation methods and found two to be particularly effective: chain-of-thought prompting, which provides a step-by-step plan for data analysis, which led to a 60% improvement in code accuracy; and self-reflection, enabling LLMs to iteratively refine their code, yielding a 38% accuracy improvement. Building on these insights, we developed a platform that integrates LLMs into the data science workflow for medical professionals. In a user study with five medical doctors, we found that while LLMs cannot fully automate coding tasks, they significantly streamline the programming process. We found that 80% of their submitted code solutions were incorporated from LLM-generated code, with up to 96% reuse in some cases. Our analysis highlights the potential of LLMs, when integrated into expert workflows, to enhance data science efficiency in clinical research.
**footnotetext: These authors contributed equally to this work.
1 Introduction
@@
In clinical research, data science plays a pivotal role in analyzing complex datasets, such as clinical trial data and real-world data (RWD), which are critical for improving patient care and advancing evidence-based medicine [1]. For example, it was reported that for a pharmaceutical company, the insights brought from RWD analysis could unlock up to \$300M values annually by optimizing clinical trial design and execution [2]. Data scientists, at the core of this process, require years of coding expertise, alongside a deep understanding of diverse medical data types, including patient clinical data and omics data, while also collaborating closely with medical professionals [3]. However, the growing demand for data science skills and the limited availability of experienced data scientists have become bottlenecks in the clinical research process [4]. Coding is central to the work of data scientists, underpinning essential tasks such as statistical modeling, data cleaning, and visualization using Python and R. Given this, exploring methods for streamlining the coding process in clinical research data science is crucial to accelerate drug development and improve patient outcomes.
@@
Code generation has been extensively explored with the advent of large language models (LLMs), which have demonstrated strong capabilities in tasks like code completion [5]. Continuous efforts have been made in developing more powerful code-specific LLMs [6, 7, 8], refining prompting strategies [9], integrating external knowledge through retrieval-augmented generation (RAG) [10, 11], and enabling LLMs’ self-reflection [12]. These advancements further lead to the LLM-based platform for software development [13] and data analysis [14]. While LLMs have been evaluated for general programming tests [5, 15, 16, 17], software engineering [18], and data analysis  [19, 20], assessments specifically targeting clinical research data science remain scarce. In medicine, recent works have introduced LLMs to automate machine learning modeling [21] and support bioinformatics tool development [22], but they do not cover broad data science tasks. Therefore, this paper seeks to build a comprehensive code generation dataset to assess to which extent the cutting-edge LLMs can automate clinical research data analysis, modeling, and visualization.
@@
Our objective was to evaluate the practical utility of LLMs in handling complex clinical research data and performing the associated data science tasks. To this end, we identified 39 clinical studies published in medical journals that were linked to patient-level datasets (Fig. 1a). We started by extracting and summarizing the analyses performed in these studies, such as patient characteristic exploration and Kaplan-Meier curves. We then developed the code necessary to reproduce these analyses and the reported results in these studies. These coding tasks, along with their reference solutions, were all manually crafted and cross-verified to ensure accuracy. The result was a collection of 293 diverse, high-quality data science tasks, covering primary tools used in Python and R, e.g., lifelines for survival analysis in Python and Bioconductor for biomedical data analysis in R. Additionally, we categorized the difficulty of these tasks into Easy, Medium, and Hard, by the number of “semantic lines” of code in the reference solutions (Fig 1d). The semantic lines metric aggregates lines of code that serve the same operation into a single unit, providing a clear measure of task complexity. A detailed overview of the dataset and its characteristics is provided in Fig. 1.
@@
Here, we rigorously evaluated the extent to which clinical research data science tasks can be automated by LLMs. We benchmarked six state-of-the-art LLMs using various methods, including chain-of-thought prompting, few-shot prompting, automatic prompting, self-reflection, and retrieval-augmented generation (Fig. 1f). Our analysis focused on both the accuracy and quality of the code generated by these models. While we found that current LLMs are not yet capable of fully automating complex clinical data science tasks, they do generate code that is highly similar to the correct final solutions. Building on this insight, we investigated the development of a platform designed to facilitate collaboration between human experts and artificial intelligence (AI) to streamline coding tasks. This platform aims to enhance the productivity of clinical researchers by integrating LLMs into established data science workflows, with a focus on user-friendliness and the reliability of the outputs. Our results demonstrated that the platform significantly improved human experts’ efficiency in executing data science tasks, highlighting the promising future of human-AI collaboration in clinical research data science.
2 Results
@@
2.1 Creating data science tasks from clinical studies
@@
We curated our testing dataset, CliniDSBench, to reflect the real-world challenges that data scientists face in clinical research. The dataset is grounded in published medical studies and linked to patient-level datasets from cBioPortal [23]. These patient datasets are diverse, each containing data from hundreds to thousands of patients, including clinical information such as demographics and survival data, lab results, and omics data like gene expression, mutations, structural variants, and copy number alterations. Unlike prior studies that mostly focus on single spreadsheets, each study in our dataset is linked to multiple spreadsheets, up to five, providing a more complex and realistic basis for evaluating data science workflows. This setup mirrors the multifaceted nature of real-world clinical research, where data scientists must integrate and analyze information from various sources to generate insights. The dataset building and evaluation framework is illustrated in Fig. 1a.
@@
We obtained the publications associated with these datasets from PubMed and reviewed them to extract the types of analyses conducted. Through this process, we identified common analyses frequently performed in clinical research, such as summarizing patient baseline characteristics, plotting Kaplan-Meier curves to assess treatment effects across groups, and creating mutational OncoPrints to highlight significant gene mutations in specific patients. This extraction process allowed us to filter and refine the initial set of studies, ensuring both diversity and comprehensiveness in the analyses covered. After this refinement, we retained 39 studies, which were used to create the final testing dataset.
@@
We designed a series of coding questions in a step-by-step format, mirroring the logical progression of analyses in the original studies, ultimately leading to the main findings. For example, a study may include exploratory data analysis, gene mutation analysis to detect abnormal mutation patterns, survival analysis to visualize patient outcomes, and statistical tests to verify significance. Correspondingly, we developed coding questions for each step, ensuring that earlier steps provide the necessary groundwork for subsequent analyses. As such, each coding task consists of five components: (1) the input question, (2) patient dataset schema description, (3) prerequisite code (called “prefix”), (4) reference solutions, and (5) test cases. This design reflects the practical setup data scientists encounter in real-world projects. In total, we manually curated 128 analysis tasks in Python and 165 in R based on the extracted analyses from 39 studies. As shown in Fig. 1c, the input questions typically consist of 50-100 words describing the task and output requirements, while the reference code solutions span more than 20 lines, sometimes exceeding 50 lines, reflecting the complexity of the tasks.
@@
We quantitatively assessed the difficulty of each coding task by calculating the number of semantic lines in the reference solutions. A semantic line aggregates multiple lines of code that contribute to the same operation, as illustrated in Fig. 1d. This approach prevents the difficulty assessment from being skewed by repetitive or tedious tasks that are fundamentally simple. The statistics for semantic lines and difficulty levels are presented in Fig. 1e. Our analysis shows that Python solutions tend to be more complex than R solutions, particularly for Medium and Hard tasks. This is largely due to R’s rich ecosystem of medical-specific packages, which often allow for more direct solutions. In contrast, Python frequently requires additional customization and manual coding to achieve similar outcomes, contributing to higher complexity in Python coding tasks. The pie charts in Fig. 1b show the libraries frequently used in the reference answers.
Figure 2: Assessment of different models and adaptation methods in automating clinical research data science tasks. a, the inputs for LLMs to generate the code and the associated evaluation process. b, the pass@5 of three LLMs with varying temperatures across difficulty levels in the Python coding dataset. c, the proportions of the reference solution code that can be drawn directly from the LLM-generated code. d and e show the pass@1 of six LLMs across difficulty levels in Python and R coding datasets, respectively.
2.2 LLMs are not yet ready for fully automated data science
@@
As illustrated in Fig. 1f, our evaluation framework is composed of three key components: models, methods, and tasks. For the first component, we selected six cutting-edge LLMs: GPT-4o [24], GPT-4o-mini [25], Sonnet [26], Opus [27], Gemini-pro [28], and Gemini-flash [28]. These models represent a diverse set of advanced generalist LLMs capable of performing code generation based on input instructions. To explore their effectiveness and potential room for improvement in clinical research data science tasks, we applied several adaptation methods: chain-of-thought [29], few-shot prompting [30], automatic prompting [31], self-reflection [12], and retrieval-augmented generation (RAG) [32].
@@
For evaluation, we assessed the combinations of models and adaptation methods across three tasks: code generation, code debugging, and human-AI collaboration. The first two tasks were used to measure the models’ accuracy in automating the coding process. We employed Pass@$k$ as the primary metric, where $k$ represents the number of attempts the model is allowed to make to solve a coding task. This metric behaves like the probability that at least one out of the $k$ attempts is correct. Specifically, we selected $k=1$ as a strict benchmark to evaluate how well LLMs can automate tasks on the first attempt, providing insight into their immediate accuracy. Additionally, we used $k=5$ as a more relaxed metric to explore the model’s potential to improve given multiple attempts, offering a broader assessment of its ability to generate correct solutions when allowed more tries.
@@
We first evaluated the immediate accuracy of LLMs in generating code solutions on their initial attempt. For each task, the LLM is provided with a raw question that describes the target task, as well as a dataset description (Fig. 2a). The dataset description includes details such as spreadsheet names, column names, and common cell values, which guide the LLM in identifying the correct spreadsheet, column, and values to work with. Additionally, the instruction section offers supplementary guidance for the LLM during code generation. We used three types of instructions: Vanilla, Manual, and Automatic. The Vanilla instruction provides minimal guidance, merely instructing the LLM to solve the task, while Manual and Automatic instructions are more detailed, either manually crafted or optimized through automatic prompt generation [31]. The generated code solutions must pass all the testing cases to be judged right.
@@
From our experiments, we found that current LLMs cannot consistently produce perfect code for clinical research data science tasks across all difficulty levels. As shown in Fig. 2d, for Python tasks, the Pass@1 scores vary significantly based on task difficulty. For Easy tasks, most LLMs achieve Pass@1 rates in the range of 0.40-0.80. However, for Medium tasks, the Pass@1 rates drop to 0.15-0.40, and for Hard tasks, they range from 0.05 to 0.15. Performance differences also exist between different LLMs, particularly within the same series. For instance, the lightweight variant GPT-4o-mini generally underperforms compared to its larger counterpart, GPT-4o, with differences in performance of up to twofold in many cases. This highlights the limitations of current LLMs, especially as task complexity increases. The trend is similar in R, where performance declines with increased task difficulty, though there is a significant difference in performance between Python and R tasks (Fig 2e).
@@
2.3 Unlocking the power of LLMs through strategic adaptations
Motivated by the varied performances of LLMs with different instruction levels, we hypothesized that tailored adaptations for LLMs in clinical research data science could lead to greater improvements. To test this, we introduced two key dimensions of adaptation: (1) enhancing LLM inference and reasoning by incorporating advanced instructions or external knowledge, and (2) employing multiple rounds of trial-and-error, allowing LLMs to iteratively correct their errors. The results of these adaptations are shown in Fig. 3.
For the first dimension of adaptation, in addition to Manual (ManualPrompt) and Automatic prompt optimization (AutoPrompt), we introduced three additional strategies: chain-of-thought (CoT), few-shot prompting (Few-shot), and retrieval-augmented generation (RAG). ManualPrompt incorporates human knowledge into the instructions, offering additional hints such as common error cases, key columns like unique patient identifiers, and specific guidance for certain analyses. AutoPrompt utilizes the dspy prompt optimizer [31], which generates prompts via an LLM and selects the best one. We optimized prompts using three studies from the training set, keeping 11 studies as the test set. RAG equips LLMs with a Google search engine, enabling them to look up package documentation, StackOverflow discussions, and clinical knowledge before generating code solutions. Few-shot prompting adds several example question-and-answer pairs from the training set to guide the model. For CoT, we enriched the instructions with step-by-step guidance, asking the LLM to follow concrete steps toward the final solution. These instructions were manually created to ensure accuracy, mimicking scenarios where proficient data scientists provide more detailed input.
@@ 2.3 Unlocking the power of LLMs through strategic adaptations
  • RAG performed similarly to Vanilla, despite incorporating external knowledge into the inputs. We hypothesize this is because GPT-4o was likely trained on a wide range of public sources, including package documentation, webpages, medical articles, and online guidelines. As a result, the additional information retrieved by RAG offered minimal benefit, as much of it was already within the model’s pre-existing knowledge. Furthermore, the retrieval process can sometimes introduce noise, embedding irrelevant or distracting context into the prompt, which negatively affects performance.
@@
We conducted further experiments to evaluate whether LLMs can solve more problems through self-reflection. The results are shown in Fig. 3c for Python tasks and Fig. 3d for R tasks. To enable self-reflection, we provided LLMs with three types of logs captured from the first attempt at executing the code solutions: (1) results from running the test cases, (2) runtime logs that capture any errors encountered during execution, and (3) additional print statements that show the values and shapes of intermediate variables. These logs were combined with the original code to help the LLM generate an explanation for the errors and propose a corrective plan, including revised code (Fig. 3a). We tested self-reflection over multiple rounds, from 1 to 5, and tracked the trend of Pass@1 performance throughout the process. From the results, we observed significant improvements through self-correction. After five rounds of correction, Pass@1 scores increased by an average of around 0.2 across all task types. For Python tasks, LLMs could solve approximately 60-80% of Easy tasks, 40-50% of Medium tasks, and 20-25% of Hard tasks after self-correction. For R tasks, LLMs achieved around 40-70% success on Easy tasks, 30-55% on Medium tasks, and 50-60% on Hard tasks. This represents a substantial improvement compared to their first attempt, demonstrating the effectiveness and potential of LLMs’ self-reflection capabilities. Notably, most of the improvement occurred within the first two rounds, with diminishing returns in later rounds, indicating that early corrections are the most impactful.
@@
We conducted an in-depth analysis of the erroneous LLM-generated solutions for Python (Fig. 3e) and R tasks (Fig. 3f), categorizing the errors into six types: Tests failure, Data misoperation, Package misuse, Instruction misfollow, Invalid syntax, and Timeout. Specifically, Data misoperation refers to errors arising from incorrect operations on the input datasets, such as selecting from non-existing columns. Package misuse includes errors from passing incorrect arguments to functions, importing incorrect packages, or calling functions without proper imports. Instruction misfollow occurs when LLMs fail to follow the provided instructions, leading to outputs that are a mixture of text and code or refusing to answer the question. A couple of example error cases are shown in Fig. 3g. Overall, most of the erroneous solutions failed the testing cases but could still be executed. The next most common errors were Data misoperation and Package misuse. After applying self-reflection, the most significant improvement came from LLMs resolving many of the Data misoperation and Package misuse errors, making the code executable, which is reflected by the increase in errors related to Tests failure. This demonstrates the utility of LLM self-correction in addressing relatively superficial errors identified through execution logs. One notable anomaly was Gemini-Pro, which encountered a high number of Instruction Misfollow errors, especially after self-reflection. This was likely due to Gemini-Pro’s strict safety policies, which caused the model to refuse to answer certain coding questions, particularly when trying to do self-correction.
@@
Upon reviewing the error cases, we found that many of the LLM-generated solutions, while imperfect, are close to correct and only require minor manual edits. To quantify how much these solutions can reduce the effort required for data science tasks, we compared the LLM-generated solutions with the reference solutions (Fig. 2c). Specifically, we conducted a difference analysis to compute the proportion of reference code that could be replicated by the LLM-generated solutions. The results show that, despite their imperfections, LLM-generated solutions are promising for streamlining data science workflows. For instance, LLMs produced code that covered approximately 60% of the reference solutions for Easy tasks, around 60% for Medium tasks in R and 40% in Python, and about 50% for Hard tasks in R and 25% in Python. It is important to note that this metric underestimates the code similarities, as AI-generated code may achieve the same function in a different way from the reference solution.
@@
2.4 Human-AI collaboration boosts productivity for data science in clinical research
@@
By far, the two critical findings from our experiments are: (1) When human experts provide more detailed, step-by-step instructions, the quality of LLM-generated code significantly improves, as demonstrated by the superior performance of Chain-of-Thought (CoT) prompting (Fig. 3b). (2) Although LLM-generated code is often imperfect, it serves as a strong starting point for human experts to refine. Evidence from Fig. 2c shows that LLM-generated code is close to the correct solution, and Fig. 3e and Fig. 3f indicate that most code executes successfully but fails only at the final testing stages. Additionally, LLM self-reflection can resolve most of the bugs. These findings highlight the potential of LLMs to assist data scientists in streamlining the coding process in clinical research.
@@
To bridge the gap in utilizing LLMs for clinical research and leveraging the insights from our experiments, we developed a platform that integrates LLMs into data science projects for clinical research. The architecture of this platform is shown in Extended Fig. 1. The platform is designed to offer an integrated interface for users to:
@@
  • Identify and debug errors in the user-provided code, with LLMs proposing solutions to improve the code.

### v2

# Can Large Language Models Replace Data Scientists in Biomedical Research?

@@ Abstract
Data science plays a critical role in biomedical research, but it requires professionals with expertise in coding and medical data analysis. Large language models (LLMs) have shown great potential in supporting medical tasks and performing well in general coding tests. However, existing evaluations fail to assess their capability in biomedical data science, particularly in handling diverse data types such as genomics and clinical datasets. To address this gap, we developed a benchmark of data science coding tasks derived from the analyses of 39 published studies. This benchmark comprises 293 coding tasks (128 in Python and 165 in R) performed on real-world TCGA-type genomics and clinical data. Our findings reveal that the vanilla prompting of LLMs yields suboptimal performances due to drawbacks in following input instructions, understanding target data, and adhering to standard analysis practices. Next, we benchmarked six cutting-edge LLMs and advanced adaptation methods, finding two methods to be particularly effective: chain-of-thought prompting, which provides a step-by-step plan for data analysis, which led to a 21% code accuracy improvement (56.6% versus 35.3%); and self-reflection, enabling LLMs to refine the buggy code iteratively, yielding an 11% code accuracy improvement (45.5% versus 34.3%). Building on these insights, we developed a platform that integrates LLMs into the data science workflow for medical professionals. In a user study with five medical professionals, we found that while LLMs cannot fully automate programming tasks, they significantly streamline the programming process. We found that 80% of their submitted code solutions were incorporated from LLM-generated code, with up to 96% reuse in some cases. Our analysis highlights the potential of LLMs to enhance data science efficiency in biomedical research when integrated into expert workflows.
Introduction
@@
In biomedical research, data science plays a pivotal role in analyzing complex datasets, such as clinical trial data and real-world data (RWD), which are critical to improving patient care and advancing evidence-based medicine [1]. For example, it was reported that for a pharmaceutical company, the insights derived from the RWD analysis could unlock up to \$300M values annually by optimizing the design and execution of clinical trials [2]. Data scientists, at the core of this process, require years of coding experience, along with a deep understanding of various types of medical data, including clinical data from patients and omics data, while also collaborating closely with medical professionals [3]. However, the growing demand for data science skills and the limited availability of experienced data scientists have become bottlenecks in the biomedical research process [4]. Coding is central to the work of data scientists, underpinning essential tasks such as statistical modeling, data cleaning, and visualization using Python and R. Given this, exploring methods for streamlining the coding process in biomedical data science is crucial to accelerate drug development and improve patient outcomes.
@@
Code generation has been extensively studied with large language models (LLMs), which have demonstrated strong capabilities in tasks like code completion [5]. Continuous efforts have been made to develop more powerful code-specific LLMs [6, 7, 8], refine prompting strategies [9], integrate external knowledge through retrieval-augmented generation (RAG) [10, 11], and enable LLMs to self-reflection [12]. These advancements lead to the LLM-based platform for software development [13] and data analysis [14]. Although LLMs have been evaluated for general programming tests [5, 15, 16, 17], software engineering [18], and data analysis  [19, 20], assessments specifically targeting biomedical data science remain scarce. In medicine, recent work has introduced LLMs to automate machine learning modeling [21] and support bioinformatic tool development [22], but they do not cover broad data science tasks or assess to what extent LLMs can collaborate with humans to complete programming tasks. Therefore, this paper seeks to build a comprehensive code generation dataset to assess to what extent cutting-edge LLMs can automate biomedical data analysis, modeling, and visualization.
@@
Our objective is to evaluate the practical utility of LLMs in handling complex biomedical data and performing associated data science tasks. To this end, we identified 39 medical and biomedical studies published in medical journals that were linked to patient-level datasets (Fig. 1a). A list of example studies and data from the linked patients can be found in Extended Fig. 1. We started by extracting and summarizing the analyses performed in these studies, such as patient characteristic exploration and Kaplan-Meier curves. We then developed the code necessary to reproduce these analyses and the results reported in these studies. These coding tasks, along with their reference solutions, were crafted and cross-verified manually to ensure accuracy. The result was a collection of 293 diverse and high-quality data science tasks, covering primary tools used in Python and R, e.g., lifelines for survival analysis in Python and Bioconductor for biomedical data analysis in R. Additionally, we categorized the difficulty of these tasks into Easy, Medium, and Hard, by the number of “semantic lines" of code in the reference solutions (Fig 1d). The semantic lines metric aggregates lines of code that serve the same operation into a single unit, providing a clear measure of task complexity. A detailed overview of the dataset and its characteristics is provided in Fig. 1.
@@
In this work, we evaluated the extent to which LLM can automate biomedical data science tasks. We benchmarked six state-of-the-art LLMs using various methods, including chain-of-thought prompting, few-shot prompting, automatic prompting, self-reflection, and retrieval-augmented generation (Fig. 1f). Our analysis focused on both the accuracy and quality of the code generated by these models. Although we found that current LLMs are not yet capable of fully automating complex biomedical data science tasks, they do generate code that is highly similar to the correct final solutions. Building on this insight, we investigated the development of an integrated platform designed to facilitate collaboration between human experts and artificial intelligence (AI) to streamline data science coding tasks. This platform aims to enhance the productivity of biomedical researchers by integrating LLMs into established data science workflows, with a focus on user-friendliness and the reliability of the outputs. Our results demonstrated that the platform significantly improved human experts’ efficiency in executing data science tasks, highlighting the promising future of human-AI collaboration in biomedical data science.
Results
@@
Creating data science tasks from medical and biomedical studies
@@
We curated our testing dataset, BioDSBench, to reflect the real-world challenges that data scientists face in biomedical research. The dataset is grounded in published medical studies and linked to patient-level datasets from cBioPortal [23]. These datasets are diverse, each containing data from hundreds to thousands of patients, including clinical information such as demographics and survival data, lab results, and omics data like gene expression, mutations, structural variants, and copy number alterations. Unlike prior studies that mostly focus on single tables, each study in our dataset is linked to multiple tables and each can be with thousands of columns, providing a more complex and realistic basis for evaluating data science workflows. This setup mirrors the multifaceted nature of real-world biomedical research, where data scientists must integrate and analyze information from various sources to generate insights. The dataset building and evaluation framework is illustrated in Fig. 1a.
@@
We obtained the publications associated with these datasets from PubMed and extracted the types of analysis in these papers. Through this process, we identified common analyses frequently performed in biomedical research, such as summarizing patient baseline characteristics, plotting Kaplan-Meier curves to assess treatment effects across groups, and creating mutational OncoPrints to highlight significant gene mutations in specific patients. This extraction process allowed us to filter and refine the initial set of studies, ensuring both diversity and comprehensiveness in the analyses covered. After this refinement, we retained 39 studies that were used to create the final testing dataset.
@@
We designed a series of coding questions in a step-by-step format, mirroring the logical progression of the analyses in the original studies, ultimately leading to the main findings. For example, a study may include exploratory data analysis, gene mutation analysis to detect abnormal mutation patterns, survival analysis to visualize patient outcomes, and statistical tests to verify significance. Correspondingly, we developed coding questions for each step, ensuring that earlier steps provide the necessary groundwork for subsequent analyses. As such, each coding task consists of five components: (1) the input question, (2) dataset schema description, (3) prerequisite code (called “prefix”), (4) reference solutions, and (5) test cases. This design reflects the practical setup that data scientists encounter in real-world projects. In total, we manually curated 128 analysis tasks in Python and 165 in R based on the extracted analyses from 39 studies. As shown in Fig. 1c, the input questions typically consist of 50-100 words describing task and output requirements, while reference code solutions span more than 20 lines, sometimes exceeding 50 lines, reflecting the complexity of the tasks.
@@
We evaluated the difficulty of each coding task by calculating the number of semantic lines in the reference solutions. A semantic line aggregates multiple lines of code that contribute to the same operation, as illustrated in Fig. 1d. This approach prevents the difficulty assessment from being skewed by repetitive or tedious tasks that are fundamentally simple. The statistics for semantic lines and difficulty levels are presented in Fig. 1e. Our analysis shows that Python solutions tend to be more complex than R solutions, particularly for Medium and Hard tasks. This is largely due to R’s rich ecosystem of medical-specific packages, which often allow for more direct solutions. In contrast, Python frequently requires additional customization and manual coding to achieve similar outcomes, contributing to higher complexity in Python coding tasks. The pie charts in Fig. 1b show the libraries frequently used in reference solutions.
LLMs are not yet ready for fully automated data science
@@
As illustrated in Fig. 1f, our evaluation framework consists of three key components: models, methods, and tasks. For the first component, we selected six cutting-edge LLMs: GPT-4o [24], GPT-4o-mini [25], Sonnet [26], Opus [27], Gemini-pro [28], and Gemini-flash [28]. These models represent a diverse set of advanced generalist LLMs capable of performing code generation based on input instructions. To explore their effectiveness and potential room for improvement in biomedical data science tasks, we applied several adaptation methods: chain-of-thought [29], few-shot prompting [30], automatic prompting [31], self-reflection [12], and retrieval-augmented generation (RAG) [32].
@@
For evaluation, we assessed the combinations of models and adaptation methods across three tasks: code generation, code debugging, and human-AI collaboration. The first two tasks were used to measure the models’ accuracy in automating the coding process. We used Pass@$k$ as the primary metric, where $k$ represents the number of attempts the model is allowed to make to solve a coding task. This metric behaves like the probability that at least one out of the $k$ attempts is correct. Specifically, we selected $k=1$ as a strict benchmark to evaluate how well LLMs can automate tasks on the first attempt, providing insight into their immediate accuracy. Moreover, employing $k=5$ served as a broader metric, enabling us to examine the model’s capability to improve with several attempts, thereby providing a more comprehensive evaluation of its potential to produce accurate solutions when given additional opportunities.
@@
We first evaluated the immediate accuracy of LLMs in generating code solutions on their initial attempt. For each task, the LLM is provided with a raw question that describes the target task, as well as a dataset description (Fig. 2a). The dataset description includes details such as table names, column names, and common cell values, which guide the LLM in identifying the correct table, column, and values to work with. Additionally, the instruction section offers supplementary guidance for the LLM during code generation. We used three types of instructions: Vanilla, Manual, and Automatic. The Vanilla instruction provides minimal guidance, merely instructing the LLM to solve the task, while Manual and Automatic instructions are more detailed, either manually crafted or optimized through automatic prompt generation [31]. The generated code solutions must pass all the testing cases to be considered right.
@@
From our experiments, we found that the vanilla adoption of LLMs cannot consistently produce perfect code for biomedical data science tasks across all difficulty levels. As shown in Fig. 2d, for Python tasks, the Pass@1 scores vary significantly based on task difficulty. For Easy tasks, most LLMs achieve Pass@1 rates in the range of 0.40-0.80. However, for Medium tasks, the Pass@1 rates drop to 0.15-0.40, and for Hard tasks, they range from 0.05 to 0.15. Performance differences also exist between different LLMs, particularly within the same series. For instance, the lightweight variant GPT-4o-mini generally underperforms compared to its larger counterpart, GPT-4o, with differences in performance of up to twofold in many cases. This highlights the limitations of current LLMs, especially as task complexity increases. The trend is similar in R, where performance declines with increased task difficulty, though there is a significant difference in performance between Python and R tasks (Fig 2e).
@@
Unlocking the power of LLMs through strategic adaptations
Motivated by the varied performances of LLMs with different instruction levels, we hypothesized that tailored adaptations for LLMs in biomedical data science could lead to greater improvements. To test this, we introduced two key dimensions of adaptation: (1) enhancing LLM inference and reasoning by incorporating advanced instructions or external knowledge, and (2) employing multiple rounds of trial-and-error, allowing LLMs to iteratively correct their errors. The results of these adaptations are shown in Fig. 3.
For the first dimension of adaptation, in addition to Manual (ManualPrompt) and Automatic prompt optimization (AutoPrompt), we introduced three additional strategies: chain-of-thought (CoT), few-shot prompting (Few-shot), and retrieval-augmented generation (RAG). ManualPrompt incorporates human knowledge into the instructions, offering additional hints such as common error cases, key columns like unique patient identifiers, and specific guidance for certain analyses. AutoPrompt utilizes the DSPy prompt optimizer [31], which generates prompts via an LLM and selects the best one. We optimized prompts for Python code generation using three studies from the training set, keeping 11 studies as the test set. RAG equips LLMs with a Google search engine, enabling them to look up package documentation, StackOverflow discussions, and clinical knowledge before generating code solutions. Few-shot prompting adds several example question-and-answer pairs from the training set to guide the model. For CoT, we enriched the instructions with step-by-step guidance, asking the LLM to follow concrete steps toward the final solution. These instructions were manually created to ensure accuracy, mimicking scenarios where proficient data scientists provide more detailed input.
@@
  • RAG performed similarly to Vanilla, despite incorporating external knowledge into the inputs. We hypothesize this is because GPT-4o was likely trained on a wide range of public sources, including package documentation, webpages, medical articles, and medical guidelines. As a result, the additional information retrieved by RAG offered minimal benefit, as much of it was already within the model’s pre-existing knowledge. Furthermore, the retrieval process can sometimes introduce noise, embedding irrelevant or distracting context into the prompt, which negatively affects performance.
@@
We conducted further experiments to evaluate whether LLMs can solve more problems through self-reflection. The results are shown in Fig. 3c for Python tasks and Fig. 3d for R tasks. To enable self-reflection, we provided LLMs with three types of logs captured from the first attempt at executing the code solutions: (1) results from running the test cases, (2) runtime logs that capture any errors encountered during execution, and (3) additional print statements that show the values and shapes of intermediate variables. These logs were combined with the original code to help the LLM generate an explanation for the errors and propose a corrective plan, including revised code (Fig. 3a). We tested self-reflection over multiple rounds, from 1 to 5, and tracked the trend of Pass@1 performance throughout the process. From the results, we observed significant improvements through self-reflection. After five rounds of correction, Pass@1 scores increased by an average of around 0.2 across all task types. For Python tasks, LLMs could solve approximately 60-80% of Easy tasks, 40-50% of Medium tasks, and 20-25% of Hard tasks after self-reflection. For R tasks, LLMs achieved around 40-70% success on Easy tasks, 30-55% on Medium tasks, and 50-60% on Hard tasks. This represents a substantial improvement compared to their first attempt, demonstrating the effectiveness and potential of LLMs’ self-reflection capabilities. Notably, most of the improvement occurred within the first two rounds, with diminishing returns in later rounds, indicating that early corrections are the most impactful.
@@
We conducted an in-depth analysis of the erroneous LLM-generated solutions for Python (Fig. 3e) and R tasks (Fig. 3f), categorizing the errors into six types: Tests failure, Data misoperation, Package misuse, Instruction misfollow, Invalid syntax, and Timeout. Specifically, Data misoperation refers to errors arising from incorrect operations on the input datasets, such as selecting from non-existing columns. Package misuse includes errors in passing incorrect arguments to functions, importing incorrect packages, or calling functions without proper imports. Instruction misfollow occurs when LLMs fail to follow the provided instructions, leading to outputs that are a mixture of text and code or refusing to answer the question. Several example error cases are shown in Fig. 3g. Overall, most of the erroneous solutions failed the testing cases but could still be executed. The next most common errors were Data misoperation and Package misuse. After applying self-reflection, the most significant improvement came from LLMs resolving many of the Data misoperation and Package misuse errors, making the code executable, which is reflected by the increase in errors related to Tests failure. This demonstrates the utility of LLM self-reflection in addressing relatively superficial errors identified through execution logs. One notable anomaly was Gemini-Pro, which encountered a high number of Instruction Misfollow errors, especially after self-reflection. This was likely due to Gemini-Pro’s strict safety policies, which caused the model to refuse to answer certain coding questions, particularly when trying to do self-reflection.
@@
Upon reviewing the error cases, we found that many of the LLM-generated solutions, while imperfect, are close to correct and only require minor manual edits. To quantify how much these LLM-generated solutions can reduce the human coding effort for data science tasks, we compared the LLM-generated solutions with the reference solutions (Fig. 2c). We conducted a difference analysis to calculate the proportion of human-validated reference code that overlaps with the solutions generated by the LLM. The results show that, despite their imperfections, LLM-generated solutions are promising for streamlining data science workflows. For instance, LLMs produced code that covered approximately 60% of the reference solutions for Easy tasks, around 60% for Medium tasks in R and 40% in Python, and about 50% for Hard tasks in R and 25% in Python. It is important to note that this metric underestimates the code similarities, as AI-generated code may achieve the same function in a different way from the reference solution.
@@
Human-AI collaboration boosts productivity for data science in biomedical research
@@
By far, the two critical findings from our experiments are: (1) When human experts provide more detailed, step-by-step instructions, the quality of LLM-generated code significantly improves, as demonstrated by the superior performance of Chain-of-Thought (CoT) prompting (Fig. 3b). (2) Although LLM-generated code is often imperfect, it serves as a strong starting point for human experts to refine. Evidence from Fig. 2c shows that LLM-generated code is close to the correct reference solution, and Fig. 3e and Fig. 3f indicate that most codes can execute successfully but fail only at the final testing stages. Additionally, LLM self-reflection can resolve most of the bugs. These findings highlight the potential of LLMs to assist data scientists in streamlining the coding process in biomedical research.
@@
To bridge the gap in utilizing LLMs for biomedical research and leveraging the insights from our experiments, we developed a platform that integrates LLMs into data science projects. The architecture of this platform is shown in Fig. 4. The platform is designed to offer an integrated interface for users to:
@@
  • Identify and bugs in the user-provided code, with LLMs proposing solutions to improve the code.